A 400.0-N sign hangs from the end of a uniform strut. The strut is 4.0 m long and

weighs 600.0 N. The strut is supported by a hinge at the wall and by a cable whose

other end is tied to the wall at a point 3.0 m above the left end of the strut. Find the

tension in the supporting cable and the force of the hinge on the strut.


1
Expert's answer
2021-12-13T16:56:09-0500

a)

"M=0,\implies"

"600\cdot 2+400\cdot 4-4T\cdot \frac 35=0,"

"T=1167~N,"

b)

"F_h=\frac45T=933~N,"

"F_v=600\cdot \frac 24=300~N,"

"F=\sqrt{F_h^2+F_v^2}=980~N."